OpenStudy (anonymous):

All real numbers at least 5 units from 8

OpenStudy (zzr0ck3r):

\[(-\infty,3]\cup[13,\infty)\]

OpenStudy (zzr0ck3r):

I think this is what you are asking, but Im not sure the question, or if this is the answer...lol

OpenStudy (anonymous):

The question was to use interval notation to define each interval. I wasn't sure if infinity classified as a "real" number.

OpenStudy (zzr0ck3r):

It's considered a extended real number, and we may write it as such. example: we may say \[\large R=(-\infty,\infty)\]

OpenStudy (anonymous):

What is the "U" looking symbol between the two notations?